The delivery trucks of Moore's Building Supply incurred maintencace costs of $2,400 during its busiest month of 20xx, in which 8,000 miles were driven collectively. During its slowest month, $1,800 in maintenance costs were incurred, reslting from 5,000 miles being driven, Using the high-low method, what maintenance cost would the company expect to incur at 10,000 miles of driving?

a.$600

b.$2,800

c.$2,000

d.$1,600
Answer #1

Variable cost per unit=[Total cost at highest level-Total cost at lowest level]/(Highest level-Lowest level)

=(2400-1800)/(8000-5000)

=$0.2/mile

Total fixed cost=2400-(8000*0.2)

=$800

Hence maintenance cost  at 10,000 miles of driving=800+(0.2*10000)

=$2800.

NOTE:Variable cost per unit and Total fixed costs do not change with change in units.